OpenStudy (anonymous):

Solve the system of linear equations below. 6x + 3y = 33 4x + y = 15 Select one of the options below as your answer: A. x = 2, y = 7 B. x = -13, y = 7 C. x = - 2 3 , y = 12 2 3 D. x = 5, y = 1

OpenStudy (texaschic101):

are you familiar with the elimination method ?

OpenStudy (anonymous):

somewhat

OpenStudy (texaschic101):

multiply the 2nd equation by -3

OpenStudy (texaschic101):

or...would you rather do substitution ?

OpenStudy (anonymous):

substitution is far easier for me

OpenStudy (texaschic101):

ok....look at the 2nd equation...can you isolate y ?

OpenStudy (anonymous):

BY adding 4x to the opposite side?

OpenStudy (anonymous):

subtracting*

OpenStudy (texaschic101):

no...subtract 4x from both sides

OpenStudy (anonymous):

so y=15-4x

OpenStudy (texaschic101):

correct....now sub 15 - 4x in for y in the 1st equation

OpenStudy (anonymous):

6x+3(15-4x)=33?

OpenStudy (texaschic101):

yes......now distribute the 3 through the parenthesis

OpenStudy (anonymous):

6x+45-12x=33. Now C.L.T right?

OpenStudy (texaschic101):

correct....combine like terms

OpenStudy (anonymous):

-6x+45=33

OpenStudy (texaschic101):

correct....now subtract 45 from both sides...or add 6x to both sides...whichever you want to do

OpenStudy (texaschic101):

but if you add 6x, you also have to subtract 33 from both sides

OpenStudy (anonymous):

-6x=-12 now divide by negative 6?

OpenStudy (texaschic101):

correct....now divide both sides by 6 keeping in mind dividing by 2 negatives gives you a positive

OpenStudy (anonymous):

so that gives me x=2 now I put that in for x on one of the equations right

OpenStudy (texaschic101):

correct

OpenStudy (anonymous):

and theen solve the same way

OpenStudy (texaschic101):

yes

OpenStudy (texaschic101):

you can then check your answers by subbing both x and y into either of the equations and if they come out equal, you have found the correct answers

OpenStudy (anonymous):

ok I put x into the first equation listed and then have gotten to 12=33-3y. What now?

OpenStudy (texaschic101):

6(2) + 3y = 33 12 + 3y = 33 how did you get that ? from here you subtract 12 from both sides

OpenStudy (anonymous):

ohhhh I subtracted 3y first

OpenStudy (texaschic101):

you are trying to get 3y on one side and everything else on the other side

OpenStudy (texaschic101):

you got it now...good job